ILAW Manufacturing company produces bulbs that last a mean of 900 hours with a standard deviation of 110 hours. what is the probability that the mean lifetime of a random sample of 15 of these bulbs is less than 850 hours?



1
Expert's answer
2021-05-18T17:45:54-0400

Let "X" be the random variable which represents the length of the life of a light bulb, in hours: "X\sim N(\mu, \sigma^2/n)"

Then



"Z={X-\mu \over \sigma/\sqrt{n}}\sim N(0,1)"

Given that "\mu=900\ hrs, \sigma=110\ hrs, n=15."


The probability that the mean lifetime of random sample of 15 of these bulbs is less than 850 hours is



"P(X<850)=P(Z<{850-900 \over 110/\sqrt{15}})\approx P(Z<-1.3176)\approx""\approx0.0938"